If the distance to the water level in a well is doubled, is the time it...

  1. If the distance to the water level in a well is doubled, is the time it takes to hear the splash twice what it was before, more than twice, or less than twice? Explain your answer.
Answer #1

If the distance to the water level in a well is doubled, the time it takes to hear the splash is less than twice what it was before.

It is so because, let initially the well depth is 5 m ,

therefore using the equation of motion, we get:

h = (1/2)gt^2

5 = (9.8/2)t^2

t = 1.01 s

Now, in case 2 , let depth of well is doubled , that is d = 10 m

10 = (9.8/2)t^2

t = 1.43 seconds

so we can easily observe from the above two considered cases that the time taken is actually less than twice.

The reason behind is this , that the object dropped into the well is under a constant acceleration (g) , thus it keeps on covering more distances in equal intervals of time as it keeps on travelling for longer and longer duration of time.

Please ask your doubts or queries in the comment section below.

Please kindly upvote if you are satisfied with the solution.

Thank you.